# Geospatial analysis of toponyms in geotagged social media posts

### v1

@@ Abstract
Place names, or toponyms, play an integral role in human representation and communication of geographic space. In particular, how people relate each toponym with particular locations in geographic space should be indicative of their spatial perception. Here, we make use of an extensive dataset of georeferenced social media posts, retreived from Twitter, to perform a statistical analysis of the geographic distribution of toponyms and uncover the relationship between toponyms and geographic space. We show that the occurrence of toponyms is characterized by spatial inhomogeneity, giving rise to patterns that are distinct from the distribution of common nouns. Using simple models, we quantify the spatial specificity of toponym distributions and identify their core-periphery structures. In particular, we find that toponyms are used with a probability that decays as a power law with distance from the geographic center of their occurrence. Our findings highlight the potential of social media data to explore linguistic patterns in geographic space, paving the way for comprehensive analyses of human spatial representations.
@@ I Introduction
When we speak or write about geographic spaces, we generally communicate them using the names of places, or toponyms. Although any point on Earth can be specified by geographic coordinates, one would rarely refer to a place in day-to-day conversation as, for example, “35.7°35.7°35.7  °N, 139.7°139.7°139.7  °E”; instead, we would usually represent it by a toponym, such as “Tokyo”. The use of toponyms reflects the way we perceive and mentally structure geographic space. Unlike geographic coordinates, which are objective and unambiguous, the area a toponym refers to is often vague and difficult to define; even for the names of administratively defined areas (such as municipalities), how people use them in colloquial settings may have only a loose correspondence with the officially demarcated boundaries [1, 2]. This however does not mean that the extent that each toponym denotes can be arbitrarily defined by individuals; since the main function of toponyms is to effectively communicate geographic information, there must be a shared consensus within the population that determines the areas they represent.
The objective of this study is to understand such a spontaneous and collective relationship between geographic space and the use of toponyms. In order to quantitatively examine this relationship, one needs to collect large scale data on the locations people associate with each toponym. The abundance of user-generated content online, especially in social media, offers a promising opportunity for this purpose. In particular, Twitter (rebranded to “X” in 2023), one of the major social media platforms, had provided free access to posts on the platform through APIs until 2023, allowing researchers to perform statistical analysis and find patterns and trends in the data. On Twitter, users could opt to attach to each of their post geolocation metadata (a geotag) that represent the geographic coordinates of the GPS location of their device. These user-annotated geotags as well as the content of the posts allow us to explore the spatial dimension of user behavior and language use. For instance, by leveraging the fact that the set of vocabulary that appears in the text of posts varies according to the whereabouts of users, studies have found that toponyms in text can be disambiguated [3] and the location of individual users can be identified [4, 5]. When aggregated at the population level, the data can be used to study dialectical variations and are of interest to sociolinguistics and linguistic geography [6, 7, 8, 9, 10].
@@
In this study, we focus on understanding the geographic distribution of geotagged posts and the toponyms they contain. Our goal is not to identify patterns of toponym usage at the individual user level, but rather to understand the collective, population-level knowledge about the relationship between toponyms and geographic locations. Similar questions have been addressed in the literature: Hollenstein and Purves [17] used user-annotated geotag data from the image hosting service Flickr to delineate city centers and neighborhoods; Hu and Janowicz [18] studied the names of points of interest (such as restaurants and shops) registered on Yelp, a user-generated local business review platform, in metropolitan areas in the United States. These studies were primarily focused on urban areas, and therefore the analysis was limited to a relatively small geographic scale. In addition, the results were rather descriptive and were not aimed at deriving general laws of toponymic occurrence. In contrast, we use a data-driven modeling approach to uncover the underlying principles governing the occurrence of toponyms of different granularities on a larger geographic scale, aiming to understand how these distributions reflect collective spatial perceptions.
We first show that the distribution of the total set of geotagged posts is characterized by a power-law density distribution, reflecting the inhomogeneity of the population distribution. We then focus on the geographic distribution of individual toponyms. We find that the occurrence of each toponym is inhomogeneously distributed in geographic space, in contrast to the distributions of common nouns and toponyms referring to a place outside the observation area. We show that such inhomogeneity cannot be explained without taking into account the spatial specificity of each toponym, and propose a simple alternative model that reproduces features of the observed data. We conclude by discussing the implications of our work.
II Material
@@
From the full sample of the $277$ million geotagged posts, we extract the subset of posts that contains each of 24 Japanese toponyms that refers to regions, prefectures, cities, and wards (special wards of Tokyo Metropolis and wards of designated cities) in Japan. The list of toponyms sampled in this work and the locations of administrative divisions they refer to can be found in Fig 1. In addition, we extracted posts that contain twelve common Japanese nouns and six Japanese toponyms that refer to places outside Japan. We call these samples for individual keywords (toponyms and nouns) as keyword subsamples. In the following, we use only the information about the number of posts tagged inside each grid cell for each sample, and disregard the content of the text or user information. Refer to Table 1 for the notation and definition of variables used in this work.
Figure 1: Places in Japan denoted by the toponyms studied in this paper. In this study, we sample 24 toponyms: the names of (A) six regions, (B) six prefectures, (C) six major cities, and (D) six wards (submetropolitan/submunicipal districts). The colored areas in each panel show the administratively defined geographic area (except for cities) denoted by each toponym. (A) The extent of each region is not uniquely defined. Here we show one of the commonly used classification of regions. (C) The colored area shows the metropolitan employment area [19], which is considered to be more representative of urban activity than administratively defined city areas. Note that Kyoto, Hiroshima, and Fukuoka are used both as the names of the cities and as the names of the prefectures of which the cities are the capitals. (D) Prefectural boundaries are shown for visual guidance.
@@ III Results
III.1 Spatial distribution of geotagged posts
@@
Figure 2: The spatial densities of geotagged posts, resident population, and employed population. (A–C) Geographic distribution of the three densities. Note that the population data are geographically limited inside Japan, while the geotagged posts are sampled in bounding box of Japan, which also includes neighboring countries. (D) Probability distribution of density (the number of geotagged posts per unit area). (E, F) Scatter plots showing the correlation between each of the population densities and the geotagged post density. Pearson and the Spearman correlation coefficients are shown below the plot. (G) Probability distributions of the three densities, each rescaled by its mean.
@@
To investigate the origin of this heterogeneous distribution, we compare the geotagged post statistics with census data for the resident population and the employed population. The employed population is defined as the number of permanent or temporary employees, self-employed individuals, contractors, and unpaid staff in family-owned businesses, whose main working site is located in each grid cell. From Figs. 2B and  2C, one can see the similarity in the spatial distribution between the densities of geotagged posts, resident population, and employed population. In fact, the geotagged post density is strongly correlated with the population densities (Figs. 2E,  2F). The correlation with the post density is stronger for the employee population density than for the resident population density, both in terms of Pearson’s correlation and Spearman’s rank correlation. This may suggest that the social media posts are made more in the daytime than at night. Figure 2G shows the probability density functions of the geotagged post density, the resident population density, and the employee population density collapse to one another when they are scaled by the average density, strongly suggesting that the uneven distribution of population gives rise to the heterogeneity in the spatial distribution of geotagged posts.
III.2 Spatial distributions of toponym subsamples
@@
We now turn to our main question: How are geotagged posts containing each toponym, i.e., each toponym subsample, spatially distributed? To address this question, we show the results for “Fukuoka”, a toponym referring to the sixth largest city in Japan as of 2019, located in the southwestern part of the country, as an illustrative example.
@@
Figure 3A shows the spatial distribution of the occurrence of “Fukuoka”, while Fig 3C shows the probability density functions of the gridwise occurrence density $\sigma_{w}$. Both figures suggest that the “Fukuoka” occurs heterogeneously across different grid cells. The same observation can be made for other toponyms, as shown in Fig S2 in Supporting Information. It is noteworthy that toponym subsamples of different sizes follow the same scaling, which implies that the spatial heterogeneity does not depend on the popularity of toponyms. On the other hand, each distribution is characterized by a single power law, which is a clear deviation from the scaling observed for the full sample.
@@
Figure 3: Occurrence pattern of “Fukuoka”. (A) Geographic distribution of density $\sigma_{w}$. (B) Geographic distribution of occurrence ratio $\phi_{w}$. (C) Probability distributions of spatial density for all geotagged posts (in gray) and for “Fukuoka” (in blue). (D) Scatter plot showing the relationship between the total number of geotagged posts $n_{\mathrm{all}}$ and the number of posts containing “Fukuoka” $n_{w}$ in each grid cell. The lines represent contours along which the density of points (kernel density estimate) on the double logarithmic scale is constant. The region inside each contour, from dark to light colors, contains 90.0%, 99.0%, and 99.9% of the data points, respectively. (E) The same scatter plot as panel D, but with points colored by the distance between the grid cell and the center $O_{w}$.
@@
While the spatial density (occurrence per unit area) of posts containing the word “Fukuoka” is high in Fukuoka and neighboring areas, it is also high in other large metropolitan areas such as Tokyo and Osaka, which are geographically distant from Fukuoka. However, this is presumably just by chance due to the large total number of posts tagged in these areas. To account for the difference in total sample size $n_{\mathrm{all}}$ in each grid cell, we normalize the toponym subsample size $n_{w}$ by $n_{\mathrm{all}}$ and obtain the fraction of occurrence of the toponym. We clearly see that posts contain word “Fukuoka” with higher probabilities in the region around the city of Fukuoka (Fig 3B).
@@
To further understand the relationship between the full sample and the toponym subsamples, we create a scatter plot for each toponym as in Fig 3D, where the horizontal axis represents $n_{\mathrm{all}}$ and the vertical axis corresponds to $n_{w}$. Apart from the general trend that $n_{w}$ increases with $n_{\mathrm{all}}$, we see that there are two distinct branches of scaling, with one increasing faster than the other. That is, $n_{w}$ increases as a function of $n_{\mathrm{all}}$ in two different ways. This two-branch scaling behavior is not specific to “Fukuoka” but is widely seen for different toponyms across various degrees of popularity and granularity (Fig S3 in Supporting Information). In contrast, keyword subsamples for common nouns such as “wallet” and toponyms that refers to places outside the observed area such as “Hawaii” do not exhibit heterogeneity in spatial distribution, and the relationship between $n_{\mathrm{all}}$ and $n_{w}$ only shows a single scaling (see Figs. 4B and 4D; see also Fig S3 for the scatter plots for all keywords sampled in this study). This suggests that the two-branch scaling is unique to the toponyms that refer to places within the observed area and is not present in other types of words. Naturally, we expect that such pattern stems from the geospatial specificity of toponym use.
@@
where $S_{w}=\{c\mid\phi_{w,c}\geq\epsilon\}$ is the set of cells with an occurrence ratio equal to or greater than $\epsilon$. This condition prevents a cell from being selected as the center merely due to the large total number of posts. Here we set $\epsilon=0.01$, i.e., toponym $w$ must appear in at least 1% of all posts tagged to a cell for the cell to be included in $S_{w}$. As we can see in Fig 3E, the grid cells that constitute the two branches are clearly distinct in terms of distance from the center of the toponym subsample. The branch with faster scaling is geographically closer to the center while the branch with slower scaling is relatively far from the center. This corroborates our hypothesis that the two-branch scaling arise from the spatial specificity of the toponym.
III.3 Location-independent model
The results in the previous section give us an intuitive understanding of the geographic distribution of toponyms. In the next two sections, we present a model-based analysis to validate our intuition and to characterize the empirical observation in a quantitative way. Specifically, we introduce a model where whether each geotagged post contains word $w$ is a Bernoulli trial with success probability $p_{w,c}$, which possibly depends on the location $c$ it is tagged to. The number $n_{w,c}$ of posts in grid cell $c$ that contains word $w$ out of $n_{\mathrm{all},c}$ total posts follows a binomial distribution:

### v2

@@ Abstract
Place names, or toponyms, play an integral role in human representation and communication of geographic space. In particular, how people relate each toponym with particular locations in geographic space should be indicative of their spatial perception. Here, we make use of an extensive dataset of georeferenced social media posts, retrieved from Twitter, to perform a statistical analysis of the geographic distribution of toponyms and uncover the relationship between toponyms and geographic space. We show that the occurrence of toponyms is characterized by spatial inhomogeneity, giving rise to patterns that are distinct from the distribution of common nouns. Using simple models, we quantify the spatial specificity of toponym distributions and identify their core-periphery structures. In particular, we find that toponyms are used with a probability that decays as a power law with distance from the geographic center of their occurrence. Our findings highlight the potential of social media data to explore linguistic patterns in geographic space, paving the way for comprehensive analyses of human spatial representations.
@@ I Introduction
When we speak or write about geographic spaces, we generally communicate them using the names of places, or toponyms. Although any point on Earth can be specified by geographic coordinates, one would rarely refer to a place in day-to-day conversation as, for example, “35.7°35.7°35.7  °N, 139.7°139.7°139.7  °E”; instead, we would usually represent it by a toponym, such as Tokyo. The use of toponyms reflects the way we perceive and mentally structure geographic space. Unlike geographic coordinates, which are objective and unambiguous, the area a toponym refers to is often vague and difficult to define; even for the names of administratively defined areas (such as municipalities), how people use them in colloquial settings may have only a loose correspondence with the officially demarcated boundaries [1, 2]. This however does not mean that the extent that each toponym denotes can be arbitrarily defined by individuals; since the main function of toponyms is to effectively communicate geographic information, there must be a shared consensus within the population that determines the areas they represent.
The objective of this study is to understand such a spontaneous and collective relationship between geographic space and the use of toponyms. In order to quantitatively examine this relationship, one needs to collect large-scale data on the locations people associate with each toponym. The abundance of user-generated content online, especially on social media, offers a promising opportunity for this purpose. In particular, Twitter (rebranded to X in 2023), one of the major social media platforms, had provided free access to posts on the platform through APIs until 2023, allowing researchers to perform statistical analysis and find patterns and trends in the data. On Twitter, users could opt to attach to each of their post geolocation metadata (a geotag) that represent the geographic coordinates of the GPS location of their device. These user-annotated geotags as well as the content of the posts allow us to explore the spatial dimension of user behavior and language use. For instance, by leveraging the fact that the set of vocabulary that appears in the text of posts varies according to the whereabouts of users, studies have found that toponyms in the text can be disambiguated [3] and the location of individual users can be identified [4, 5], although there are limitations to this approach at high spatial resolution [6]. When aggregated at the population level, the data can be used to study dialectal variation and language evolution, making it of interest to sociolinguistics and linguistic geography [7, 8, 9, 10, 11, 12].
@@
In this study, we focus on understanding the geographic distribution of geotagged posts and the toponyms they contain. Our goal is to understand the collective, population-level knowledge about the relationship between toponyms and geographic locations rather than to identify patterns of toponym usage at the individual user level. Similar questions have been addressed in the literature: Hollenstein and Purves [19] used user-annotated geotag data from the image hosting service Flickr to delineate city centers and neighborhoods; Hu and Janowicz [20] studied the names of points of interest (such as restaurants and shops) registered on Yelp, a user-generated local business review platform, in metropolitan areas in the United States. These studies were primarily focused on urban areas, and therefore the analysis was limited to a relatively small geographic scale. In addition, the results were rather descriptive and were not aimed at deriving general laws of toponymic occurrence. In contrast, we use a data-driven modeling approach to uncover the underlying principles governing the occurrence of toponyms of different granularities on a larger geographic scale, aiming to understand how these distributions reflect collective spatial cognition.
The rest of the paper is structured as follows. We start by introducing the geotagged Twitter dataset and how we collect, preprocess, and subsample it in the Data section. In the Results section, we first present the inhomogeneity of the geographic distribution of geotagged posts, and compare it to the population distributions. We then focus on the geographic distribution of individual toponyms. In the second subsection, we observe that the occurrence of each domestic toponym follows a characteristic pattern, hinting at its spatial specificity. To formalize this observation, we introduce a class of models called binomial models. In the third subsection, we use the simplest instance of this model to quantify the spatial specificity of each toponym. In the last subsection, we show that another variant of the binomial model, which we call the core-periphery model, reproduces the essential elements of toponym occurrence patterns despite its simplicity. In the Discussion and Conclusions section, we discuss the implications of our work.
II Data
We collected 395 268 777395268777395\,268\,777395 268 777 geotagged Twitter posts from the Twitter API. These posts are annotated with coordinates within the bounding box of Japan (latitude between 20.43°20.43°20.43  °N and 45.56°45.56°45.56  °N and longitude between 122.93°122.93°122.93  °E and 153.99°153.99°153.99  °E) during the period from 1 February 2012 to 30 September 2018. Note that this bounding box also includes South and North Korea as well as parts of China and Russia. The data collection was conducted in accordance with Twitter’s Terms of Service, which allow researchers to analyze and publish findings based on Twitter data, but prohibit the redistribution of raw data, such as the text or geotags of individual posts.
Data collection was followed by several preprocessing stages. First, we excluded posts made via location check-in services (e.g., Foursquare) as well as those generated by automated bots or manipulative users. Posts made through Foursquare are in specific text formats, such as “I’m at [the name of a place/point of interest]” or “[post text] (@ [the name of a place/point of interest])”. When a user checks in to a place, they select the name of the place/point of interest, which may include toponyms, from a list of nearby places provided by Foursquare. However, users do not have independent control over the geographic coordinates tagged to the post; these coordinates are automatically determined by Foursquare’s location database based on the name of the place chosen. Although these posts are created by personal users, they do not represent an organic association between toponyms and geographic coordinates. Including them in the analysis would bias the findings of this study.
[…]
@@ II Data
From the full sample of the $276$ million geotagged posts, we extract the subset of posts that contains each of 24 Japanese toponyms that refer to regions, prefectures, cities, and wards (special wards of Tokyo Metropolis and wards of designated cities) in Japan. The list of toponyms sampled in this work and the administrative areas they refer to can be found in Fig 1.
[…]
In addition to these domestic toponyms, we extracted posts that contain twelve common Japanese nouns and six Japanese toponyms that refer to places outside Japan to construct reference datasets. We refer to the samples for individual keywords (toponyms and nouns) as keyword subsamples. In the following, we only use the information about the number of posts tagged inside each grid cell for each sample, and disregard the content of the text or user information. Refer to Table 1 for the notation and definition of variables used in this work.
Figure 1: Places in Japan denoted by the toponyms studied in this paper. In this study, we sample 24 toponyms: the names of (A) six regions, (B) six prefectures, (C) six major cities, and (D) six wards (submetropolitan/submunicipal districts). The colored areas in each panel show the administratively defined geographic area (except for cities) denoted by each toponym. (A) The extent of each region is not uniquely defined. Here we show one of the commonly used classifications of regions. (C) The colored area shows the metropolitan employment area [23, 24], which is considered to be more representative of urban activity than administratively defined city areas. Note that Kyoto, Hiroshima, and Fukuoka are used both as the names of the cities and as the names of the prefectures of which the cities are the capitals. (D) Prefectural boundaries are shown for visual guidance. Maps made with Natural Earth (https://www.naturalearthdata.com/).
@@ III Results
Spatial distribution of geotagged posts
@@
Figure 2: The spatial densities of geotagged posts, resident population, and employed population. (A–C) Geographic distribution of the three densities. Note that the population data are geographically limited inside Japan, while the geotagged posts are sampled in the bounding box of Japan, which also includes neighboring countries. Maps made with Natural Earth (https://www.naturalearthdata.com/). (D) Probability distribution of density (the number of geotagged posts per unit area). (E, F) Scatter plots showing the correlation between each of the population densities and the geotagged post density. Pearson and the Spearman correlation coefficients are shown below the plot. (G) Probability distributions of the three densities, each rescaled by its mean.
@@
To investigate the origin of this heterogeneous distribution, we compare the geotagged post statistics with census data for the resident population in 2015 [25] and the employed population in 2016 [26]. The employed population is defined as the number of permanent or temporary employees, self-employed individuals, contractors, and unpaid staff in family-owned businesses, whose main working site is located in each grid cell. From Figs. 2B and  2C, one can see the similarity in the spatial distribution between the densities of geotagged posts, resident population, and employed population. In fact, the geotagged post density is strongly correlated with the population densities (Figs. 2E,  2F). The correlation with the post density is stronger for the employee population density than for the resident population density, both in terms of Pearson’s correlation and Spearman’s rank correlation. This may suggest that social media posts are made more in the daytime than at night. Figure 2G shows the probability density functions of the geotagged post density, the resident population density, and the employee population density collapse to one another when they are scaled by the average density, strongly suggesting that the uneven distribution of population gives rise to the heterogeneity in the spatial distribution of geotagged posts.
III.1 Spatial distributions of toponym subsamples
@@
We now turn to our main question: How are geotagged posts containing each toponym, i.e., each toponym subsample, spatially distributed? To address this question, we show the results for Fukuoka, a toponym referring to the sixth largest city in Japan as of 2019, located in the southwestern part of the country, as an illustrative example.
@@
Figure 3A shows the spatial distribution of the occurrence of Fukuoka, while Fig 3C shows the probability density functions of the gridwise occurrence density $\sigma_{w}$. Both figures suggest that the Fukuoka occurs heterogeneously across different grid cells. The same observation can be made for other toponyms, as shown in Fig S1 in Supporting Information. It is noteworthy that toponym subsamples of different sizes follow the same scaling, which implies that the spatial heterogeneity does not depend on the popularity of toponyms. On the other hand, each distribution is characterized by a single power law, which is a clear deviation from the scaling observed for the full sample.
@@
Figure 3: Occurrence pattern of Fukuoka. (A) Geographic distribution of density $\sigma_{w}$. (B) Geographic distribution of occurrence ratio $\phi_{w}$. (C) Probability distributions of spatial density for all geotagged posts (in gray) and for Fukuoka (in blue). (D) Scatter plot showing the relationship between the total number of geotagged posts $n_{\mathrm{all}}$ and the number of posts containing Fukuoka $n_{w}$ in each grid cell. The lines represent contours along which the density of points (kernel density estimate) on the double logarithmic scale is constant. The region inside each contour, from dark to light colors, contains 90.0%, 99.0%, and 99.9% of the data points, respectively. (E) The same scatter plot as panel D, but with points colored by the distance between the grid cell and the center $O_{w}$. Maps made with Natural Earth (https://www.naturalearthdata.com/).
@@
While the spatial density (occurrence per unit area) of posts containing the word Fukuoka is high in Fukuoka and neighboring areas, it is also high in other large metropolitan areas such as Tokyo and Osaka, which are geographically distant from Fukuoka. However, this is presumably just by chance due to the large total number of posts tagged in these areas. To account for the difference in total sample size $n_{\mathrm{all}}$ in each grid cell, we normalize the toponym subsample size $n_{w}$ by $n_{\mathrm{all}}$ and obtain the fraction of occurrence of the toponym. We clearly see that posts contain the word Fukuoka with higher probabilities in the region around the city of Fukuoka (Fig 3B).
@@
To further understand the relationship between the full sample and the toponym subsamples, we create a scatter plot for each toponym as in Fig 3D, where the horizontal axis represents $n_{\mathrm{all}}$ and the vertical axis corresponds to $n_{w}$. Apart from the general trend that $n_{w}$ increases with $n_{\mathrm{all}}$, we see that there are two distinct branches of scaling, with one increasing faster than the other. That is, $n_{w}$ increases as a function of $n_{\mathrm{all}}$ in two different ways. This two-branch scaling behavior is not specific to Fukuoka but is widely seen for different toponyms across various degrees of popularity and granularity (Fig S2 in Supporting Information). In contrast, keyword subsamples for common nouns such as wallet and toponyms that refer to places outside the observed area such as Hawaii do not exhibit heterogeneity in spatial distribution, and the relationship between $n_{\mathrm{all}}$ and $n_{w}$ only shows a single scaling (see Figs. 4B and 4D; see also S2 for the scatter plots for all keywords sampled in this study). This suggests that the two-branch scaling is unique to the toponyms that refer to places within the observed area and is not present in other types of words. Naturally, we expect that such a pattern stems from the geospatial specificity of toponym use.
@@
where $S_{w}=\{c\mid\phi_{w,c}\geq\epsilon\}$ is the set of cells with an occurrence ratio equal to or greater than $\epsilon$. This condition prevents a cell from being selected as the center merely due to the large total number of posts. Here we set $\epsilon=0.01$, i.e., toponym $w$ must appear in at least 1% of all posts tagged to a cell for the cell to be included in $S_{w}$. As we can see in Fig 3E, the grid cells that constitute the two branches are clearly distinct in terms of distance from the center of the toponym subsample. The branch with faster scaling is geographically closer to the center while the branch with slower scaling is relatively far from the center. This corroborates our hypothesis that the two-branch scaling arises from the spatial specificity of the toponym.
III.2 Location-independent model
The results in the previous section give us an intuitive understanding of the geographic distribution of toponyms. In the next two sections, we present a model-based analysis to validate our intuition and to characterize the empirical observation in a quantitative way. Specifically, we introduce a model where each post tagged to location $c$ contains word $w$ with probability $p_{w,c}$ that may depend on $c$. We assume that the occurrence of word $w$ in each post is an independent event, i.e., a Bernoulli trial. The number $n_{w,c}$ of posts in grid cell $c$ that contains word $w$ out of $n_{\mathrm{all},c}$ total posts follows a binomial distribution: